### v2

@@ Abstract
Hypercomplex neural network, algebra, tensor
pacs:
MSC Classification]15A69, 15-04
@@ 1 Introduction
The fast progress in applications of Artificial Neural Networks (NN) promotes new directions of research and generalizations. This involves advanced mathematical concepts such as group theory [21], differential geometry [5, 6], or topological methods in data analysis [7].
@@
There are different algebraic extensions. One of these paths follows Algebraic Neural Networks [16], where the additional endomorphism operations on data are performed. The other algebra-geometry direction is the neural networks based on Geometric Algebra/Clifford algebra [4, 19], starting from the most prominent example of Quaternions [9, 12]. Recently, Parametrized Hypercomplex Neural Networks were invented [11] for convolutional layers. They can learn optimal hypercomplex algebra adjusted to data exploring optimized Kronecker product. However, some applications need hyperalgebra, or even more general algebra parameters, as (fixed) hyperparameters. In such a way we can optimize algebra structure at the metalevel.
In this paper, we discuss implementation in which we change classical real algebra computations into various hypercomplex or even more general algebra computations by including multiplication in the algebra as a tensorial operation. Hypercomplex NN approach is not new, e.g., [3, 9]. However, interest in this direction is revived due to better complexity properties in such areas as image processing [20] or time series analysis [13]. In these contributions the Open Source code [20] for specific four-dimensional hypercomplex algebras were used.
The implementation explained in this article significantly expands the ideas from [20, 9] for arbitrary algebras, including hypercomplex ones. The algorithms described here agree with the NN presented in [20, 9] when limited to 4-dimensional hypercomplex algebras. The 4-dimensional hyperalgebras are useful in encoding RGB color data in images. Moreover, implementation of [20] was obtained by constructing an additional multiplication matrix from the multiplication table for the hypercomplex algebra. This is an additional step in setting up a neural network. Our approach permits us to omit this complexity and generalize to arbitrary algebras since we input only the multiplication table of the algebra as it can be found in standard references, e.g., [2]. Moreover, the limitation to 4-dimensional algebras can be relaxed. This is important for processing data that naturally encodes in $n$-tuples ($n>0$, integer) that can be encoded as a single element of an $n$-dimensional algebra. This is a crucial contribution from the theoretical treatment of the general algebraic approach to hypercomplex neural networks.
@@
  • provide a general algorithm for computations within the hypercomplex dense layer that generalizes ideas of [20, 9] for arbitrary algebras of arbitrary algebra dimensions,
  • provide general algorithms for 1-, 2-, and 3-dimensional hypercomplex convolutional layer computations that generalizes ideas of [20, 9] for arbitrary algebras of any algebra dimensions.
@@ 2.1 Tensors
Corollary 1.
When the algebra is commutative, then $A(x,y)=A(y,x)$ or $A_{ij}^{k}=A_{ji}^{k}$.
@@
Example 5.
For complex numbers $\mathbb{C}$, $V$ has two dimensional base $\{e_{0},e_{1}\}$ with multiplication matrix given in [16].
@@ 3 Neural Networks Architectures
We want to stress out that the proposed algorithms are generalizations of algorithms from [20, 9] from $4$-dimensional algebras to arbitrary dimensional algebras using tensor operations that are optimized for tensor operations from packages like TensorFlow or PyTorch. They are implemented in Hypercomplex Keras Python library [22] and described elsewhere [15]. The implementation was constructed to agree for $4$-dimensional algebras with the results from [20].
@@ 3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions; cf. [10].

### v3

@@ Abstract
Tensor, Hypercomplex-valued Neural Network, Hypercomplex Algebra, TensorFlow, PyTorch
journal: ArXiv
[…]
[inst1]organization=Faculty of Computer Science and Telecommunications, Cracow University of Technology, addressline=Warszawska 24, city=Cracow, postcode=PL31-155, country=Poland
@@ 1 Introduction
The fast progress in applications of Artificial Neural Networks (NN) promotes new directions of research and generalizations. This involves advanced mathematical concepts such as group theory [24], differential geometry [6], [7], or topological methods in data analysis [8].
@@
There are different algebraic extensions. One of these paths follows Algebraic Neural Networks [18], where the additional endomorphism operations on data are performed. The other algebra-geometry direction is the neural networks based on Geometric Algebra/Clifford algebra [5], [21], starting from the most prominent example of Quaternions [11], [14]. Recently, Parametrized Hypercomplex Neural Networks were invented [13] for convolutional layers. They can learn optimal hypercomplex algebra adjusted to data exploring optimized Kronecker product. However, some applications need hyperalgebra, or even more general algebra parameters, as (fixed) hyperparameters. In such a way we can optimize algebra structure at the metalevel.
In this paper, we discuss implementation in which we change classical real algebra computations into various hypercomplex or even more general algebra computations by including multiplication in the algebra as a tensorial operation. Hypercomplex NN approach is not new, e.g., [3], [11]. However, interest in this direction is revived due to better complexity properties in such areas as image processing [23] or time series analysis [15]. In these contributions the Open Source code [23] for specific four-dimensional hypercomplex algebras were used.
The implementation explained in this article significantly expands the ideas from [23], [11] for arbitrary algebras, including hypercomplex ones. The algorithms described here agree with the NN presented in [23], [11] when limited to 4-dimensional hypercomplex algebras. The 4-dimensional hyperalgebras are useful in encoding RGB color data in images. Moreover, implementation of [23] was obtained by constructing an additional multiplication matrix from the multiplication table for the hypercomplex algebra. This is an additional step in setting up a neural network. Our approach permits us to omit this complexity and generalize to arbitrary algebras since we input only the multiplication table of the algebra as it can be found in standard references, e.g., [2]. Moreover, the limitation to 4-dimensional algebras can be relaxed. This is important for processing data that naturally encodes in $n$-tuples ($n>0$, integer) that can be encoded as a single element of an $n$-dimensional algebra. This is a crucial contribution from the theoretical treatment of the general algebraic approach to hypercomplex neural networks.
@@
  • 2.
    provide a general algorithm for computations within the hypercomplex dense layer that generalizes ideas of [23], [11] for arbitrary algebras of arbitrary algebra dimensions,
  • 3.
    provide general algorithms for 1-, 2-, and 3-dimensional hypercomplex convolutional layer computations that generalizes ideas of [23], [11] for arbitrary algebras of any algebra dimensions.
@@ 2.1 Tensors
Corollary When the algebra is commutative, then $A(x,y)=A(y,x)$ or $A_{ij}^{k}=A_{ji}^{k}$.
@@
Example For complex numbers $\mathbb{C}$, $V$ has two dimensional base $\{e_{0},e_{1}\}$ with multiplication matrix given in [16].
@@ 3 Neural Networks Architectures
We want to stress out that the proposed algorithms are generalizations of algorithms from [23], [11] from $4$-dimensional algebras to arbitrary dimensional algebras using tensor operations that are optimized for tensor operations from packages like TensorFlow or PyTorch. They are implemented in Hypercomplex Keras Python library [25] and described elsewhere [17]. The implementation was constructed to agree for $4$-dimensional algebras with the results from [23].
@@ 3.1 Hypercomplex dense layer
We start with a description of the dense hypercomplex layer. It is a general-purpose layer that operates on the data with additional dimensionality/axis, which is a multiple of algebra dimensions. We assume that the input data are of dimension $b\times\underbrace{al\times in}_{n}$, where $b$ is the batch size, $al$ the algebra size, and $in$ the positive integer multiplier. The last two numbers determine the input data size $n$. The input tensor in abstract index notation is $X=X_{i_{b}Rs(i_{al},i_{in})}$, where $i_{b}$ batch index, $i_{al}$ is the algebra index, $i_{in}$ is the multiplicity index of algebra dimension. Moreover, we use learning parameters (weights/kernel) $K=K_{i_{al}i_{in}i_{u}}$, where $i_{u}$ is the index over units/neurons. The bias $b=b_{Rs(i_{al},i_{u})}$ is used if needed. Kernel and bias are usually initialized with numbers taken from specific distributions (cf. [12]).
@@ 3.2 Hypercomplex convolutional layer
4 Tensor version of universal approximation theorem for single layer of hypercomplex perceptrons
In this section we give the proof of Universal Approximation Theorem in the tensor framework outlined above. It presents how this approach provides more streamlined approach. We will follow [3]. For alternative vector-matrix approach see [22], which is also based on [3].
In this section we assume that the algebra $(V,+,\cdot)$ is non-degenerate, i.e.,
Definition 1.
An algebra $(V,+,\cdot)$ with multiplication given by the tensor $A$ is non-degenerate if for $x\in V$
$\begin{array}[]{cccc}A(x,y)=0,&\forall y\in V&\Rightarrow&x=0,\end{array}$ (18)
and
$\begin{array}[]{cccc}A(y,x)=0,&\forall y\in V&\Rightarrow&x=0.\end{array}$ (19)
The crucial part is played by the Riesz’s representation theorem that in standard form is [4]
Lemma 1 (Riesz’s representation lemma).
For a Hilbert space $H$ over field $\mathbb{F}$ (e.g., $\mathbb{R}$ or $\mathbb{C}$) with inner product <_|_$<\_|\_>:X\times X\rightarrow\mathbb{F}$ every continous linear functional $\phi\in H^{*}$ there is a vector $v_{\phi}\in H$ such that
$\phi(x)=<v_{\phi}|x>,$ (20)
for all $x\in H$.
When we take $H=\mathbb{R}^{N}$, for integer $N>0$, then $<x|y>:=x^{T}y=\sum_{i=1}^{n}x_{i}y_{i}$. Moreover, if $dim(H)<\infty$, then the linear functional is automatically continuous.
In order to amend the lemma for vectors that entries are hypercomplex numbers from $V$, we define projection on components of hypercomplex numbers $\pi_{k}(\sum_{i=0}^{n}x^{i}e_{i})=x^{k}$ for $0\leq k\leq n$. We then prove the generalization of Lemma 5.5.1 from [3]
Lemma 2 (Riesz’s representation lemma for nondegenerate algebra).
Let $\phi:V^{N}\rightarrow\mathbb{R}$ be a linear functional, where $V$ is a nondegenerate algebra. Then there exists, uniquely determined for $\phi$, the set of vectors $\{v_{i}\}_{i=0}^{n}$ for $v_{i}\in V^{N}$ such that for every $x\in V^{N}$ we have
$\phi(x)=\pi_{i}(<v_{i}|x>)$ (21)
for all $i\in\{0,\ldots,n\}$, and $<v|x>=\sum_{k=1}^{N}v_{k}\cdot x_{k}=v^{T}x$.
Proof.
We fix $0\leq i\leq n$ and we have $\phi(x)=\pi_{i}(<v_{i}|x>)$. Define a morphism $\Phi:V^{n}\rightarrow(V^{n})^{*}$ such that $\Phi(v_{i})=\phi=\pi_{i}\circ<v_{i}|\_>$. We prove that $\Phi$ defines an isomorphism.
One can note that $\Phi$ is linear and the image and preimage have the same dimension. Therefore, in order to prove that $\Phi$ is an isomorphism, we prove that $\Phi$ is injective, i.e., if $\Phi(v_{i})(x)=\pi_{i}(<v_{i}|x>)=\phi(x)=0$ for all $x\in V^{N}$, then $v_{i}=0$.
Selecting base vectors $E_{k}=[0,\ldots,0,\underbrace{1}_{k},0,\ldots,0]\in V^{N}$ for $1\leq k\leq N$ we define vectors $y_{pq}=e_{p}E_{q}$ where index $0\leq p\leq n$ runs over algebra base of $V$, and $q$ runs over components of $V^{N}$. The vectors $y_{pq}$ form a base of vectors with components from the algebra $V$. Assuming that
$0=\phi(y_{pq})=\pi_{i}(<v_{i}|y_{pq}>),$ (22)
for all $p$ and $q$, and taking into account that $V$ algebra is nondegenerate, we have that all components of $v_{i}$ must vanish, so $v_{i}=0$. As $i$ is arbitrary, this is true for all $0\leq i\leq n$ so the lemma is proved. ∎
We can now prove the Universal Approximation Theorem for hypercomplex-valued vectors. For the proof based on vector-matrix notation see [22].
Theorem 1.
Let $X\subset V^{N}$ be a compact subset and let $g:X\rightarrow V^{N}$ be a continuous function. Then for all $\varepsilon>0$ there exist some coefficients:
$\alpha_{1},\ldots,\alpha_{p}\in\mathbb{R},$ (23)
some vectors:
$v_{1},\ldots,v_{p}\in V^{N}$ (24)
and some elements of algebra:
$\Theta_{1},\ldots,\Theta_{p}\in V,$ (25)
such that:
$sup_{x\in X}||g(x)-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})||<\varepsilon,$ (26)
for $f=\sum_{i=1}^{n}\sigma\circ\pi_{i}$, where $\sigma$ is the sigmoid function, and the norm is a vector space norm on an algebra $V$ treated as a vector space.
In other words, the set of functions $S=\sum_{k=1}^{p}\alpha_{k}f(<v_{k}|x>+\Theta_{k})$ for parameters: $\alpha_{k}$, $v_{k}$ and $\Theta_{k}$ are dense in $C^{0}(X,V)$ continuous functions from $X$ to $v$ with the supremum norm over components of $V$ treated as a vector space.
Proof.
The general idea of the proof is as follows:
  1. 1.
    Decompose the function $g$ into components of the algebra $V$. The component is a real-valued function.
  2. 2.
    Use classical Universal Approximation Theorem for the component. This hides all Functional Analysis parts of the proof to already proven the classical theorem [9].
  3. 3.
    Compose components approximations into the approximation of whole function using supremum norm for estimate approximation quality.
To begin with, consider a function $g:X\rightarrow V$ that can be decomposed into components of algebra $V$ as
$g=\sum_{i=0}^{n}e_{i}\otimes\pi_{i}\circ g.$ (27)
Then each component $\pi_{i}\circ g\in C^{0}(X,\mathbb{R})$ can be approximated by the Cybenko’s Universal Approximation Theorem [9]: there are real coefficients $a_{1}^{i},\ldots,a_{p}^{i}$, real-valued linear functionals $\phi_{1}^{i},\ldots,\phi_{p}^{i}$ on $\mathbb{R}^{nN}=V^{n}$ and some real parameters $\theta_{1}^{i},\ldots,\theta_{p}^{i}$, where the index depends on the component $i$, i.e., $p=p(i)$, such that
$\left|\pi_{i}(g(x))-\sum_{k=1}^{p}a_{k}^{i}\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})\right|<\varepsilon$ (28)
for all $x\in X$.
From the Lemma 2 we have that $\phi_{k}^{i}=\pi_{i}\circ<v_{k}^{i}|\_>$ for a fixed vector $v_{k}^{i}\in V^{N}$.
Consider now $\Theta_{k}^{i}(\lambda^{i})=\lambda^{i}e_{0}+\ldots+\lambda^{i}e_{i-1}+\theta_{k}^{i}e_{i}+\lambda^{i}e_{i+1}+\ldots+\lambda^{i}e_{n}$. We have that
$lim_{\lambda^{i}\rightarrow-\infty}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))=\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})\otimes e_{i}$ (29)
uniformly. This results from the properties of $\sigma$ function ($lim_{\lambda\rightarrow-\infty}\sigma(\pi_{l}(<v_{k}|x>)+\lambda)=0$), which guarantees that all components apart the $i$-th one vanish in the limit. Therefore, we can select such $\lambda^{i}$ that
$sup_{x\in X}||\sum_{k=1}^{p}\alpha_{k}\sigma(\phi_{k}^{i}(x)+\theta_{k}^{i})e_{i}-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))||<\varepsilon,$ (30)
so we get
$sup_{x\in X}||\pi_{i}(g(x))e_{i}-\sum_{k=1}^{p}\alpha_{k}f(<v_{k}^{i}|x>+\Theta_{k}^{i}(\lambda^{i}))||<2\varepsilon.$ (31)
Since it works for all components $\pi_{i}\circ g$ of $g$, so it works also for supremum of all components. Therefore $g$ can be arbitrarily approximated by $f$. The proof is complete. ∎
The above proof that mimic the proof of Theorem 5.5.1 from [3], and it delegates functional analytic parts of the proof in application of Cybenko’s Universal Approximation Theorem [9]. One can generalize the proof by replacing the sigmoid function $\sigma$ by the discriminatory sigmoidal functions [9].